40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 279 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  2835   Fri Apr 23 18:30:49 2010 Aidan, Jenne, KojiSummaryGreen LockingGreen means GO!

Jenne, Koji and I assembled the Covesion Oven today, inserted a PPKTP crystal from Raicol, aligned the crystal to a 50mW focus and
got some green beam coming out.

Covesion Oven assembly

The oven contains a brass clip that can clamp a crystal up to 10mm wide x 0.5mm high x 40mm long (according to the instructions). According to the correspondence from Covesion the clip can accomodate a crystal up to 1.5mm high. Our crystal is 1mm x 1mm x 30mm.

  1. We removed the brass springs from the clip - see Koji's photos
  2. We placed the Raicol PPKTP crystal (#725) into the clamp with the long polished surfaces facing out to the sides and the roughened surfaces facing up and down.
  3. We balanced the 10mm x 40mm x 1mm glass plate on top of the crystal.
  4. We replaced the brass springs in the top of the clip but only tightened the screws a couple of turns so they wouldn't fall out.
  5. Very carefully and slowly, I tightened the screws a few turns in a star-shaped order to distribute the pressure evenly across the glass top
  6. Each time I tightened all eight screws, I jiggled each of the four springs to see if there was any compression in them
  7. Once the springs started to show signs of compression I stopped tightening them and tested the stability of the glass plate - a reasonable amount of pressure was required to move the plate - about the same amount required to push a SR560 across an optical table with your index finger.
  8. We loosely attached the lid and moved the oven to the table

Alignment of the crystal to the focus

The oven was mounted on a 4-axis Newport translation stage. We plonked the assembly onto the table, removed the lid and adjusted the rough position so that a focus of the 1064nm beam, from a 100mm lens, was positioned near the center of the crystal - then it was clamped down to the table. From here we adjusted the alignment of the stage, using an IR card and a viewer to guide us, until we eventually saw some green beam coming out. We were all very excited by this! We optimized the alignment as best we could using the IR card and then we replaced the lid on the oven. At this point the temperature of the PPKTP was around 26.5C and the green beam coming out look quite dim. We turned the oven up to around 36 degC and observed the beam getting much brighter and we approached the optimum phase-matching condition.

We haven't done anyway quantitative measurements yet but we were pleased with how easy this first stage was.

 

[Edit by Koji] More photos are on Picasa album

Attachment 1: IMG_2405.jpg
IMG_2405.jpg
Attachment 2: IMG_2417.jpg
IMG_2417.jpg
  2834   Thu Apr 22 21:42:24 2010 AlbertoUpdatePSLInnolight 2W Vertical Beam Profile

 

 What kind of fit did you use? How are the uncertainties in the parameters obtained?

  2833   Thu Apr 22 20:28:40 2010 JenneUpdatePEMSeismometer Noise Characteristics

Quote from elog 2237, 11 Nov 2009:

The attached plot shows the spectra of the 3 Z axes of the 3 seismometers we have (this data is from ~20Aug2009, when the Ranger was in the Z orientation) in Magenta, Cyan and Green, and the noise of each of the sensors in Red, Blue and Black.  The noise curves were extracted from the spectra using the Huddle Test / 3 Corner Hat method.  The Blue and Black traces which are just a few points are estimates of the noise from other spectra.  The Blue points come from the Guralp Spec Sheet, and the Black comes from the noise test that Rana and I did the other day with the Ranger (elog 2223).  

I'm not really happy with the black spectra - it looks way too high.  I'm still investigating to see if this is a problem with my calibration/method....

 So, as it turns out (surprise), I'm a spaz and forgot a 2*pi when calibrating the Guralp noise spectra from the spec sheet.  I noticed this when redoing the Huddle Test, and comparing my Spec Sheet Guralp noise with Rana's, which he shows in elog 2689.  When going from m/s^2, the units in the spec sheet, I just tilted the line by a factor of frequency.  Koji pointed out that I needed a factor of 2*pi*f.  That moves the Guralp spec line in the plot in elog 2237 (to which this entry is a reply) down by ~6, so that my measured noise is not, in fact, below the spec.  This makes things much more right with the world.

In other news, I redid the Huddle analysis of the 2 Guralp seismometers, ala Rana's elog 2689. The difference is now we are on the granite slab, with soft rubber feet between the floor and the granite.  We have not yet cut holes in the linoleum (which we'll do so that we're sitting directly on the 40m's slab).

Rana> this seems horrible. Its like there's a monster in there at 6-7 Hz! Either the seismos are not centered or the rubber balls are bad or Steve is dancing on the granite slab again.

 

Attachment 1: Gur1_Gur2_noise.png
Gur1_Gur2_noise.png
  2832   Thu Apr 22 15:44:34 2010 josephbUpdateComputers 

I updated the default FILTER.adl file located in /home/controls/cds/advLigo/src/epics/util/ on megatron.  I moved the yellow ! button up slightly, and fixed the time string in the upper right.

Attachment 1: New_example_CDS_filter.png
New_example_CDS_filter.png
  2831   Thu Apr 22 09:03:54 2010 steveOmnistructurePEMaircondition can not be turned off

Koji and I wanted to turn off the IFO-room AC so the wind would not blow on MC1-3. We could not. The switches were probably bypassed when the power transformer was replaced at the last scheduled power outage.

  2830   Wed Apr 21 23:35:37 2010 KojiUpdatePSLInnolight 2W Vertical Beam Profile

Good fit. I assumed sqrt(x) is a typo of sqrt(2).

Quote:

Koji and Kevin measured the vertical beam profile of the Innolight 2W laser at one point.

This data was taken with the laser crystal temperature at 25.04°C and the injection current at 2.092A.

The distance from the razor blade to the flat black face on the front of the laser was 13.2cm.

The data was fit to the function y(x)=a*erf(sqrt(x)*(x-x0)/w)+b with the following results.

Reduced chi squared = 14.07

x0 = (1.964 +-  0.002) mm

w = (0.216 +- 0.004) mm

a = (3.39 +- 0.03) V

b = (3.46 +- 0.03) V

 

  2829   Wed Apr 21 22:11:48 2010 ranaUpdatePSLInnolight 2W Vertical Beam Profile

Back in Gainesville in 1997, I learned how to do this using the chopper wheel. We had to make the assumption that the wheel's blade was moving horizontally during the time of the chop.

One advantage is that the repetitive slices reduces the random errors by a lot - you can trigger the scope and average. Another advantage is that you can download the average scope trace using USB, floppy, or ethernet instead of pencil and paper.

But, I never analyzed it in enough detail to see if there was some kind of nasty systematic error.

  2828   Wed Apr 21 21:56:27 2010 KevinUpdatePSLInnolight 2W Vertical Beam Profile

Koji and Kevin measured the vertical beam profile of the Innolight 2W laser at one point.

This data was taken with the laser crystal temperature at 25.04°C and the injection current at 2.092A.

The distance from the razor blade to the flat black face on the front of the laser was 13.2cm.

The data was fit to the function y(x)=a*erf(sqrt(x)*(x-x0)/w)+b with the following results.

Reduced chi squared = 14.07

x0 = (1.964 +- 0.002) mm

w  = (0.216 +- 0.004) mm

a  = (3.39  +- 0.03) V

b  = (3.46  +- 0.03) V

Attachment 1: bp2.jpg
bp2.jpg
Attachment 2: bp2.dat
razor height (mm)   Voltage (V)
2.75    6.89
2.50    6.90
2.30    6.89
2.25    6.89
2.20    6.75
2.15    6.47
2.13    6.20
2.10    6.05
2.07    5.88
... 17 more lines ...
  2827   Wed Apr 21 21:46:53 2010 KojiUpdateIOOMC spot diagnosis

Zach and Koji

We measured uncalibrated angle-to-length coupling using tdssine and tdsdmd.
We made a simple shell script to measure the a2l coupling.

Details:

- Opened the IMC/OMC light door.

- Saw the large misalignment mostly in pitch. Aligned using MC2 and MC3.

- Locked the MC in the low power mode. (script/MC/mcloopson AND MC length gain 0.3->1.0)

- Further aligned MC2/3. We got the transmission of 0.16, reflection of 0.2

- Tried to detect angle-to-length coupling so that we get the diagnosis of the spot positions.

- Tried to use ezcademod. Failed. They seems excite the mirror  but returned NaN.

- We used tdssine and tdsdmd instead. Succeeded.

- We made simple shell script to measure the a2l coupling. It is so far located users/koji/100421/MCspot

- We blocked the beam on the PSL table. We closed the chamber and left.

 

  2826   Wed Apr 21 16:48:38 2010 josephbUpdateCDSHardware update

Alberto and myself went to downs and acquired the 3rd 4x processor (Dual core, so 8x cores total) computer.  We also retrieved 6 BIO interface boards (blue front thin boxes), 4 DAC interface boards, and 1 ADC interface boards.  The tops have not been put on yet, but we have the tops and a set of screws for them.  For the moment, these things have been placed behind the 1Y6 rack and under the table behind the 1Y5 rack

.irwin.jpg

The 6 BIO boards have LIGO travelers associated with them: SN LIGO-S1000217 through SN LIGO-S1000222.

  2825   Wed Apr 21 15:53:57 2010 JenneUpdatePEMSeismometers now on the granite slab

The 3 seismometers are now on the granite slab.  The Ranger is now aligned with the Xarm (perpendicular to the Mode Cleaner) since that's the only way all 3 would fit on the slab.

Attachment 1: P4210048.JPG
P4210048.JPG
  2824   Wed Apr 21 11:32:31 2010 josephbUpdateCDS40m CDS hardware update and software requests

This is mostly a reminder to myself about what I discussed with Jay and Alex this morning.

The big black IO chassis are "almost" done.  Except for the missing parts.  We have 2 Dolphin, 1 Large and 1 Small I/O Chassis due to us.  One Dolphin is effectively done and is sitting in the test stand.  However, 2 are missing timing boards, and 3 are missing the boards necessary for the connection to the computer.  The parts were ordered a long time ago, but its possible they were "sucked to one of the sites" by Rolf (remember this is according to Jay).  They need to either track them down in Downs (possibly they're floating around and were just confused by the recent move), get them sent back from the sites, or order new ones (I was told by one person that the place they order from them notoriously takes a long time, sometimes up to 6 weeks.  I don't know if this is exaggeration or not...).  Other than the missing parts, they still need to wire up the fans and install new momentary power switches (apparently the Dolphin boards want momentary on/off buttons).  Otherwise, they're done.

We are due another CPU, just need to figure out which one it was in the test stand.

6 more BIO boards are done.  When I went over the plans with Jay, we realized we needed 7 more, not 6, so they're putting another one together.  Some ADC/DAC interface boards are done.  I promised to do another count here, to determine how many we have, how many we need, and then report that back to Jay before I steal the ones which are complete.  Unfortunately, he did not have a new drawing for the ASC/vertex wiring, so we don't have a solid count of stuff needed for them.  I'll be taking a look at the old drawings and also looking at what we physically have.

I did get Jay to place the new LSC wiring diagram into the DCC (which apparently the old one never was put in or we simply couldn't find it).  Its located at: https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?docid=10985

I talked briefly with Alex, reminded him of feature requests and added a new one:

1) Single part representing a matrix of filter banks

2) Automatic generation of Simulated shared memory locations and an overall on/off switch for ADC/DACs

3) Individual excitation and test point pieces (as opposed to having to use a full filter bank).  He says these already exist, so when I do the CVS checkout, I'll see if they work.

 

I also asked where the adl default files lived, and he pointed me at ~/cds/advLigo/src/epics/util/

In that directory are FILTER.adl, GDS_TP.adl, MONITOR.adl.  Those are the templates.  We also discovered the timing signal at some point was changed from something like SYS-DCU_ID to FEC-DCU_ID, so I basically just need to modify the .adl files to fix the time stamp channel as well.  I basically need to do a CVS checkout, put the fixes in, then commit back to the CVS.  Hopefully I can do that sometime today.

I also brought over 9 Contec DO-32L-PE boards, which are PCIe isolated digital output boards which do into the IO chassis.  These have been placed above the 2 new computers, behind the 1Y6 rack.

 

  2823   Wed Apr 21 10:09:23 2010 kiwamuUpdateGreen Lockingwaist positon of Gaussian beam in PPKTP crystals

Theoretically the waist position of a Gaussian beam (1064) in our PPKTP crystal differs by ~6.7 mm from that of the incident Gaussian beam.

So far I have neglected such position change of the beam waist in optical layouts because it is tiny compared with the entire optical path.

But from the point of view of practical experiments, it is better to think about it.

In fact the result suggests the rough positioning of our PPKTP crystals;

we should put our PPKTP crystal so that the center of the crystal is 6.7 mm far from the waist of a Gaussian beam in free space.


(How to)

The calculation is very very simple.

The waist position of a Gaussian beam propagating in a dielectric material should change by a factor of n, where n is the refractive index of the material.

In our case, PPKTP has  n=1.8, so that the waist position from the surface of the crystal becomes longer by n.

Now remember the fact that the maximum conversion efficiency can be achieved if the waist locates at exact center of a crystal.

Therefore the waist position in the crystal should be satisfied this relation; z*n=15 mm, where z is the waist position of the incident beam from the surface and 15 mm is half length of our crystal.

Then we can find z must be ~8.3 mm, which is 6.7 mm shorter than the position in crystal.

The attached figure shows the relation clearly. Note that the waist radius doesn't change.

Attachment 1: focal_positin_edit.png
focal_positin_edit.png
  2822   Tue Apr 20 20:15:37 2010 KevinUpdatePSLInnolight 2W Output Power vs Injection Current

Koji and Kevin measured the output power vs injection current for the Innolight 2W laser.

The threshold current is 0.75 A.

 

The following data was taken with the laser crystal temperature at 25.04ºC (dial setting: 0.12).

Injection Current (A) Dial Setting Output Power (mW)
0.000 0.0 1.2
0.744 3.66 1.1
0.753 3.72 4.6
0.851 4.22 102
0.954 4.74 219
1.051 5.22 355
1.151 5.71 512
1.249 6.18 692
1.350 6.64 901
1.451 7.08 1118
1.556 7.52 1352
1.654 7.92 1546
1.761 8.32 1720
1.853 8.67 1855
1.959 9.05 1989
2.098 9.50 2146

 

Attachment 1: PvsI_2W.jpg
PvsI_2W.jpg
  2821   Tue Apr 20 19:37:02 2010 KojiUpdateGreen Locking1W NPRO output profile

Beautiful fitting.

Quote:

EDIT: I used an IFIT (inverse fast idiot transform) to change the x-axis of the plot from Hz to m. I think xlabel('Frequency [Hz]') is in my muscle memory now..

I have redone the beam fit, this time omitting the M2, which I believe was superfluous. I have made the requested changes to the plot, save for the error analysis, which I am still trying to work out (the function I used for the least squares fit does not work out standard error in fit parameters). I will figure out a way to do this and amend the plot to have error bars.

 
profile_fit_4_19_10.png

 

  2820   Tue Apr 20 18:02:22 2010 JenneUpdateCOCNew SRM and PRM Hung

[Jenne, Steve]

We removed the old SRM and PRM from their cages, and are temporarily storing them in the rings which we use to hold the optics while baking.  Steve will work on a way to store them more permanently.

We then hung the new SRM (SRMU03) and new PRM (SRMU04) in the cages.  We were careful not to break the wires, so the heights will not have changed from the old heights.

The optics have not been balanced yet.  That will hopefully happen later this week.

  2819   Tue Apr 20 13:37:36 2010 JenneUpdateGreen Locking1W NPRO output profile

Quote:

I have redone the beam fit, this time omitting the M2, which I believe was superfluous. I have made the requested changes to the plot, save for the error analysis, which I am still trying to work out (the function I used for the least squares fit does not work out standard error in fit parameters). I will figure out a way to do this and amend the plot to have error bars.

 
profile_fit_4_19_10.png

 Are you sure about your x-axis label? 

  2818   Tue Apr 20 13:02:14 2010 ZachUpdateGreen Locking1W NPRO output profile

EDIT: I used an IFIT (inverse fast idiot transform) to change the x-axis of the plot from Hz to m. I think xlabel('Frequency [Hz]') is in my muscle memory now..

I have redone the beam fit, this time omitting the M2, which I believe was superfluous. I have made the requested changes to the plot, save for the error analysis, which I am still trying to work out (the function I used for the least squares fit does not work out standard error in fit parameters). I will figure out a way to do this and amend the plot to have error bars.

 
profile_fit_4_19_10.png
  2817   Tue Apr 20 13:00:52 2010 ZachUpdateelogelog restarted

 I restarted the elog with the restart script as it was down.

  2816   Tue Apr 20 11:14:31 2010 AidanUpdateGreen LockingRaicol crystals arrived and we investigated them

 

 Here is Crystal 724 polished side 2 with all photos along the length stitched together

  2815   Tue Apr 20 10:55:10 2010 steveBureaucracySAFETYKevin Kuns received safety training

The 40m's new undergrad Kevin Kuns was introduced to 40m safety hazards. He is new and needs guidance as specially with 2W laser work.

Peter King will train him on Friday to LIGO-laser standard.

 

  2814   Tue Apr 20 09:15:15 2010 steveSummarySAFETYannual safety audit
Attachment 1: safa.PDF
safa.PDF
  2813   Tue Apr 20 08:00:52 2010 steveUpdateSUSETMY damping restored

ETMY sus damping was restored

  2812   Tue Apr 20 07:48:42 2010 steveUpdatePSLion pump HV turned on

We found ref-cavity HV was off yesterday afternoon. It was turned back on.

  2811   Tue Apr 20 00:32:30 2010 JenneUpdatePEMGuralp Breakout Box put back

Quote:

I pulled the Guralp breakout box from the rack, and it's sitting on the EE bench here.  The game plan is to check out the Gur2X channel.  

Rana and Steve have been investigating, and found that the X channel has been funky (which has been known for ~a month or two) when the seismometer has been plugged in, and also when the seismometers have been unplugged, but the box is left on.  The funkyness goes away when the box is turned off.  Since it's not there when the box is off, it seems that it's not a problem with the cable from the box to the ADC, or in the ADC channel.  Since it is there when the box is on, but the seismometer is unplugged, it's clear that it's probably in the box itself.  

Preliminarily, I've connected a set of BNC clipdoodles to the input testpoints, and another set to the output.  They're both connected to a 'scope (which is on it's battery so it's not connected to any Ground), and when I tap on the circuit board the input trace is totally unchanged, but the output trace goes kind of crazy, and gets more fuzzy, and picks up a DC offset.  Koji is concerned that some of the big capacitors may have an iffy connection to the board. 

Investigations will continue Monday morning. 

 The Guralp Box appears to be back in working order.  It's reinstalled with the 2 seismometers plugged in.

In order:

* Koji suggested retouching the through-board solder joints on the broken channel (EW2 = Gur2X) with a bit of solder to ensure the connections were good.  Check.

* "C7", one of the giant 1uF capacitors on each channel is totally bypassed, and since that was one of the original suspects, Rana removed the (possibly) offending capacitor from EW2.

* Rana and I isolated the craziness to the final differential output stage.  We tried each of the testpoints after the individual gain / filter stages, and found that the signals were all fine, until after the output stage.

* I started to remove the resistors in the output stage (with the plan to go through the resistors, capacitors, and even the amplifier chip if neccessary), and noticed that 2 of the 1k resistors came off too easily, as if they were just barely connected in the first place.  After replacing only the 4 1k resistors, the craziness seemed to be gone.  I poked and gently bent the board, but the output wouldn't go crazy.  I declared victory.

* I reinstalled the box in its normal spot, and put Gur2 (which had been out by the bench for use as a test signal) back next to the other seismometers.  We are in nominal condition, and should be able to do a huddle test this week.

 

I looked at the time traces of all the seismometer channels, and they all looked good.  I'll put a spectra in in the morning....I'm too impatient to wait around for the low frequency FFTs.

 

Attached are the before and after pictures of the output stage of EW2 / Gur2X.  The "before" is the one with the OUT+ and OUT- words upsidedown.  The "after" picture has them right side up.

Attachment 1: Gur2X_before_20Apr2010.jpg
Gur2X_before_20Apr2010.jpg
Attachment 2: Gur2X_after_replace_100k_resistors_20Apr2010.jpg
Gur2X_after_replace_100k_resistors_20Apr2010.jpg
  2810   Mon Apr 19 16:31:42 2010 KevinUpdatePSLInnolight 2W Laser

Koji and Kevin

We unpacked the Innolight 2W laser, took an inventory, and scanned the operations manual.

[Edit by KA]

The scanned PDFs are placed on the following wiki page

http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/PSL

We will measure the P-I curve, the mode profile, frequency actuator responses, and so on.

  2809   Mon Apr 19 16:27:13 2010 AidanUpdateGreen LockingRaicol crystals arrived and we investigated them

Jenne, Koji and I opened up the package from Raicol and examined the crystals under the microscope. The results were mixed and are summarized below. There are quite a few scratches and there is residue on some of the polished sides. There is a large chip in one and there appear to be gaps or bands in the AR coatings on the sides.

There are two albums on Picassa

1. The package is opened ...

2. The crystals under the microscope.

 

Crystal Summary
724 Chip in the corner of one end face, Otherwise end faces look clean. Large scratch on one polished side.
725 End faces look good. Moderate scratch on one polished face. Residue on one polished face.
726 Tiny dot on one end face, otherwise look okay. Large bands in one polished face. Moderate scratch on polished face
727 Large, but shallow chip on one polished face. End faces look clean. Bands in one of the polished faces.

 

  2808   Mon Apr 19 13:23:03 2010 josephbConfigurationComputersyum update fixed on control room machines

I went to Ottavia, and tried running yum update.  It was having dependancy issues with mjpegtools, which was a rpmforge provided package.  In order to get it to update, I moved the rpmforge priority above (a lower number) that of epel ( epel -> 20 from 10, rpmforge -> 10 to 20).  This resolved the problem and the updates proceeded (all 434 of them). yum update on Ottavia now reports nothing needs to be done.

I went to Rosalba and found rpmfusion repositories enabled.  The only one of the 3 repositories in each file enabled was the first one.

I then added priority listing to all the repositories on Rosalba.  I set CentOS-Base and related to priority=1.  I set CentOS-Media.repo priority to 1 (although it is disabled - just to head off future problems). I set all epel related to priorities to 20. I set all rpmforge related priorities to 10.  I set all rpmfusion related priorities to 30, and left the first repo in rpmfusion-free-updates and rpmfusion-nonfree-updates were enabled.  All other rpmfusion testing repositories were disabled by me.

I then had to by hand downgrade expat to expat-1.95.8-8.3.el5_4.2.x86_64 (the rpmforge version).  I also removed and reinstalled x264.x86_64.  Lastly I removed and reinstalled lyx. yum update was then run and completed successfully.

I installed yum-priorities on Allegra and made all CentOS-Base repositories priority 1. I similarly made the still disabled CentOS-Media priority 1.  I made all epel related repos priority 20.  I made all lscsoft repos priority=50 (not sure why its on Allegra and none of the other ones).  I made all rpmforge priorities 10.  I then ran "yum update" which updated 416 packages.

 

So basically all the Centos control room machines are now using the following order for repositories:

CentOS-Base > rpmforge > epel > (rpmfusion - rosalba only) > lscsoft (allegra only)

I'm not sure if rpmfusion and lscsoft are necessary, but I've left them for now.  This should mean "yum update" will have far fewer problems in the future.

 

  2807   Mon Apr 19 11:31:04 2010 AidanUpdateGreen Locking1W NPRO output profile

Quote:

 Koji asked me to take a profile of the output of the 1W NPRO that will be used for green locking. I used the razor-scan method, plotting the voltage output of a PD vs the position of the razor across the beam, both vertically and horizontally. This was done at 6 points along the beam path out of the laser box.

I determined the beam spot size at each point by doing a least-squares fit on the plots above in Matlab (using w as one of the fitting parameters) to the cumulative distribution functions (error functions) they should approximate.

I then did another least-squares fit, fitting the above "measured" beam profiles to the gaussian form for w vs z. Below is a summary.

It seems reasonable, though I know that M2 < 1 is fishy, as it implies less divergence than ideal for that waist size. Also, like Koji feared, the waist is inside the box and thus the scan is almost entirely in the linear regime.

profile_fit_4_17_10.png

There is a clearly a difference in the divergence angle of the x and y beams - maybe 10-20%. Since the measurements are outside the Rayleigh range and approximately in the linear regime, the slope of the divergence in this plot should be inversely proportional to the waists - meaning the x- and y- waist sizes should differ by about 10-20%. You should check your fitting program for the waist.

 

  2806   Mon Apr 19 07:38:07 2010 ranaHowToElectronicsRepair and Calibration of SR560: s/n 59650

Frank noticed that this particular SR560 had an offset on the output which was unzeroable by the usual method of tuning the trim pot accessible through the front panel.

I tried to zero the offset using the trimpots inside, but it became clear that the offset was due to a damaged FET, so Steve ordered ~20 of the (now obsolete*) NPD5564.

I replaced this part and adjusted the offsets and balanced the CMRR of the differential inputs mostly according to the manual (p. 17). There are a few notes that should be added to the procedure:

  1. It can sometimes be that the gain proscribed by the manual is too high and saturates the output for large offsets. If that's the case, simply lower the gain, trim the offset, then return the gain to the specified value and trim again.
  2. The limit in trimming the offset is the stick slip resolution in the trim pot. This can potentially leave the whole preamp in an acoustically sensitive state. I tapped the pots with a screwdriver after tuning to make sure it was in more of a 'sticky' rather than 'slippy' region of the knob. A better design would allow for more filtering of the pot.
  3. In the CMRR tuning procedure it says to 'null sine wave output' but it should really say 'null the sine wave component at the drive frequency'. The best CMRR tuning uses a 1 kHz drive and leaves a residual 2 kHz signal due to the distortion imbalance (of the FETs I think).
  4. The CMRR tuning upsets the DC offset trim and vice versa. The best tuning is gotten by iterating slightly (go back and forth once or twice between the offset and CMRR tuning procedures).

It looks like its working fine now. Steve's ordering some IF3602 (low-noise, balanced FET pair from Interfet) to see if we can drop the SR560's input noise to the sub-nV level.

Noise measured with the input terminated with a BNC short (not 50 Ohms) G=100, DC coupled, low-noise mode:

Input referred noise (nV/rHz)
f e_n

0.1

200
1 44
10 8
100 5
1000 5
10000 4
  2805   Mon Apr 19 05:54:50 2010 ranaConfigurationPSLRC Temperature Servo Turned OFF temporarily

In order to measure the transfer function of the RC cavity's foam, I've turned off the servo so that the room temperature noise can excite it.

The attached plot shows a step response test from 2 weeks ago. Servo is nominally still working fine.

Attachment 1: Untitled.png
Untitled.png
  2804   Sat Apr 17 18:30:12 2010 ZachUpdateGreen Locking1W NPRO output profile

NOTE: This measurement is wrong and only remains for documentation purposes.

Koji asked me to take a profile of the output of the 1W NPRO that will be used for green locking. I used the razor-scan method, plotting the voltage output of a PD vs the position of the razor across the beam, both vertically and horizontally. This was done at 6 points along the beam path out of the laser box.

I determined the beam spot size at each point by doing a least-squares fit on the plots above in Matlab (using w as one of the fitting parameters) to the cumulative distribution functions (error functions) they should approximate.

I then did another least-squares fit, fitting the above "measured" beam profiles to the gaussian form for w vs z. Below is a summary.

It seems reasonable, though I know that M2 < 1 is fishy, as it implies less divergence than ideal for that waist size. Also, like Koji feared, the waist is inside the box and thus the scan is almost entirely in the linear regime.

profile_fit_4_17_10.png

  2803   Fri Apr 16 17:46:54 2010 KojiUpdateVACPeeting mirrors aligned

Steve and Koji

We aligned the peeping mirrors to look at the surface of the ITMs.
They had been misligned as we move the positions of the ITMs, but now they are fine.

  2802   Fri Apr 16 17:26:23 2010 JenneUpdatePEMGuralp Breakout Box pulled

I pulled the Guralp breakout box from the rack, and it's sitting on the EE bench here.  The game plan is to check out the Gur2X channel.  

Rana and Steve have been investigating, and found that the X channel has been funky (which has been known for ~a month or two) when the seismometer has been plugged in, and also when the seismometers have been unplugged, but the box is left on.  The funkyness goes away when the box is turned off.  Since it's not there when the box is off, it seems that it's not a problem with the cable from the box to the ADC, or in the ADC channel.  Since it is there when the box is on, but the seismometer is unplugged, it's clear that it's probably in the box itself.  

Preliminarily, I've connected a set of BNC clipdoodles to the input testpoints, and another set to the output.  They're both connected to a 'scope (which is on it's battery so it's not connected to any Ground), and when I tap on the circuit board the input trace is totally unchanged, but the output trace goes kind of crazy, and gets more fuzzy, and picks up a DC offset.  Koji is concerned that some of the big capacitors may have an iffy connection to the board. 

Investigations will continue Monday morning. 

  2801   Thu Apr 15 14:47:28 2010 steveUpdateElectronics25MHZ oscillation of HP4195A

The 1979 vintage RF spectrum analyzer HP4195A  sn2904J01587 shipped out  for repair today to http://www.avalontest.com

It has a 25 MHZ oscillation when you go  below 150 MHZ in your sweep....atm1 with the larger amplitude shows this 25 MHZ

Atm2 is displaying  full sweep-sign scans from 1 to 500 MHZ.....here one can clearly see the three segment of the scan:

1, large amplitude 25 MHZ oscillation dominating the spectrum up to 150 MHZ

2, the mid section from 150 MHZ  to 300 MHZ with medium size amplitude is normal

3, from 300 MHZ to 500 MHZ the amplitude is decreasing.......showing the disadvantage of using a 300 MHZ oscilloscope

 

 

 

Attachment 1: P1060246.JPG
P1060246.JPG
Attachment 2: P1060249.JPG
P1060249.JPG
  2800   Tue Apr 13 20:02:02 2010 KojiUpdateSUSBS chamber opened, PRM/SRM SOS removed from the table

Bob, Steve, and Koji

We opened North heavy door of the BS chamber in the afternoon.

In the evening, Koji worked on the PRM/SRM removal.

- Cleaned up the OPLEV mirrors to create some spaces near the door.

- Clamped PRM/SRM.

- Removed OSEMs. Made a record of the OSEMs. The record is on the wiki (http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/Suspensions)

- Found the SOSs are quite easy to remove from the table as they are shorter than the MOSs.

- Put a new Al sheet on a wagon. Put the SOSs on it. Wrapped them by the Al foils.

- Carried it to the clean room. They are on the right flow bench. Confirmed the wires are still fine.

- Closed up the chamber putting a light door.

Attachment 1: IMG_2384.jpg
IMG_2384.jpg
  2799   Tue Apr 13 19:53:06 2010 MottUpdateGreen LockingPZT response for the innolight and lightwave

 

 I redid the PZT Phase Modulation measurement out to 5 MHz for both the Innolight and the Lightwave.  The previous measurement stopped at 2MHz, and we wanted to see if there were any sweet spots above 2MHz.  I also reduced the sweep bandwidth and increased the source amplitude at high frequency to reduce the noise (the Lighwave measurement, especially, was noise dominated above 1MHz).  I also plotted the ratio of PM/AM in rad/RIN, since this is the ultimate criterion on which we want to make a determination.

It looks like there is nothing extremely useful above 2MHz for either laser.  There are several candidates for the lightwave at about 140 kHz and again at about 1.4 MHz.  The most compelling peak, however, is in the innolight at 216 kHz, where the peak is about 2.3e5 rad/RIN.

Below about 30kHz, the loop suppresses the measurement, so one should focus on the region above there.

Attachment 1: Innolight_PM.png
Innolight_PM.png
Attachment 2: Innolight_AM_PM.png
Innolight_AM_PM.png
Attachment 3: Innolight_PM_AM_Ratio.png
Innolight_PM_AM_Ratio.png
Attachment 4: Lightwave_PM.png
Lightwave_PM.png
Attachment 5: Lightwave_AM_PM.png
Lightwave_AM_PM.png
Attachment 6: Lightwave_PM_AM_Ratio.png
Lightwave_PM_AM_Ratio.png
  2798   Tue Apr 13 12:49:35 2010 josephbUpdateComputersY end simulated plant progress

Quote:

Currently, the y end plant is yep.mdl.  In order to compile it properly (for the moment at least) requires running the normal makefile, then commenting out the line in the makefile which does the parsing of the mdl, and rerunning after modifying the /cds/advLigo/src/fe/yep/yep.c file.

The modifications to the yep.c file are to change the six lines that look like:

"plant_mux[0] = plant_gndx"  into lines that look like "plant_mux[0] = plant_delayx".  You also have to add initialization of the plant_delayx type variables to zero in the if(feInt) section, near where plant_gndx is set to zero.

This is necessary to get the position feedback within the plant model to work properly.

 

#NOTE by Koji

CAUTION:
This entry means that Makefile was modified not to parse the mdl file.
This affects making any of the models on megatron.

 To prevent this confusion in the future, at Koji's suggestion I've created a Makefile.no_parse_mdl in /home/controls/cds/advLIGO on megatron.  The normal makefile is the original one (with correct parsing now).  So the correct procedure is:

1) "make yep"

2) Modify yep.c code

3) "make -f Makefile.no_parse_mdl yep"

  2797   Tue Apr 13 12:39:51 2010 Aidan, MottSummaryGreen LockingTemperature sweep of the Innolight: df/dT ~ 3.3GHz/K

Please put those numbers onto wiki somewhere at the green page or laser characterization page.

Quote:

Quote:

The beams from the Innolight and Lightwave NPROs were both incident on a 1GHZ New Focus PD. Mott and I swept the temperature of the Lightwave and tracked the change in frequency of the beatnote between the two. The Innolight temperature was set to 39.61C although the actual temperature was reported to be 39.62C.

Freq. vs temperature is plotted below in the attached PDF. The slope is 2.8GHz/K.

The data is in the attached MATLAB file.

 Same thing for the Innolight Mephisto.

Not unexpected values with dn/dT around 11E-6 K^-1 and coefficient of thermal expansion = 8E-6 K^-1 and a laser resonator length of order 10cm.

 

  2796   Mon Apr 12 22:51:31 2010 KojiUpdateSUSITMX installed and aligned

Koji

ITMX was aligned with regard to the 40m green oplev.
Now both cavities are aligned.

Next thing we are going to do is to remove PRM and SRM towers.

As well as the oplev construction for ITMs.

We anticipate the drift of the stack. So we need to revisit the alignment again.

Some tools and the level gauge were removed from the table.

Picture of the ETMX - reflection from the ITMX is hitting the mirror and Jamie's windmill.

 IMG_2381.jpg

 


0. The suspension tower had been placed on the table close to the door.

1. Brought the OSEMs from the clean room. Connected the satellite box to the ITMX suspension.

2. Went into the chamber. Leveled the table.

3. Released the mirror from the clamp. Put and adjust the OSEMs.

- Note that the side OSEM is located to the south side of the tower
  so that we can still touch it after the placement of the TT suspension at the north side of the SOS tower.

4. Clamped the mirror. Moved the SOS tower according to the CAD layout.

5. Leveled the table again.

6. Released the mirror again and adjusted the OSEMs.

7. Turned on the end green laser pointer.

- The spot was slightly upside and left of the mirror. Adjusted it so that the spot is at the center.

8. Align ITMX in Pitch

- The spot was hitting the tube. Moved the pitch bias such that the beam get horizontal.

9. Align ITMX in Yaw

- Moved the SOS tower such that the approximate spot is on the ETMX. If I hit the right spot I could see the tube get grown green because of the huge scatter.

10. Adjusted the OSEMs again and check the alignment again. Repeated this process 2~3 times.

- Bias values at  the end of the work: Pitch 0.7800 / Yaw 0.270

11. Close up the chamber

- Remove the level gauge. Some of the screws are still in the Al ship in the chamber.

- Close the light door.

  2795   Mon Apr 12 22:44:30 2010 KojiUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Data looks perfect ... but the fitting was wrong.

Vn = Vdn + Z * sqrt( 2 e Idc ) ==> WRONG!!!

Dark noise and shot noise are not correlated. You need to take a quadratic sum!!!

Vn^2 = Vdn^2 + Z^2 *(2 e Idc)

And I was confused whether you need 2 in the sqrt, or not. Can you explain it?
Note that you are looking at the raw RF output of the PD and not using the demodulated output... 

Also you should be able to fit Vdn. You should put your dark noise measurement at 10nA or 100nA and then make the fitting.

Quote:

 Here's another measurement of the noise of the REFL11 PD.

This time I made the fit constraining the Dark Noise. I realized that it didn't make much sense leaving it as a free coefficient: the dark noise is what it is.

2010-04-09_REFL11NoiseMeasurements.png

Result: the transimpedance of REFL11at 11 MHz is about 4000 Ohm.

Note:
This time, more properly, I refer to the transimpedance as the ratio between Vout @11Mhz / Photocurrent. In past entries I improperly called transimpedance the impedance of the circuit which resonates with the photodiode.

 

  2794   Mon Apr 12 20:48:51 2010 Aidan, MottSummaryGreen LockingTemperature sweep of the Innolight: df/dT ~ 3.3GHz/K

Quote:

The beams from the Innolight and Lightwave NPROs were both incident on a 1GHZ New Focus PD. Mott and I swept the temperature of the Lightwave and tracked the change in frequency of the beatnote between the two. The Innolight temperature was set to 39.61C although the actual temperature was reported to be 39.62C.

Freq. vs temperature is plotted below in the attached PDF. The slope is 2.8GHz/K.

The data is in the attached MATLAB file.

 Same thing for the Innolight Mephisto.

Not unexpected values with dn/dT around 11E-6 K^-1 and coefficient of thermal expansion = 8E-6 K^-1 and a laser resonator length of order 10cm.

Attachment 1: Innolight_temp_sweep.pdf
Innolight_temp_sweep.pdf
Attachment 2: Innolight_Temp.m
% plot the data from the Innolight Temperature sweep

% Innolight temperature

InnTemp = [0.60
    .59
    .56
    .52
    .65] + 39;

... 25 more lines ...
  2793   Mon Apr 12 19:50:30 2010 AidanSummaryGreen LockingTemperature sweep of the Lightwave: df/dT = 2.8GHz/K

The beams from the Innolight and Lightwave NPROs were both incident on a 1GHZ New Focus PD. Mott and I swept the temperature of the Lightwave and tracked the change in frequency of the beatnote between the two. The Innolight temperature was set to 39.61C although the actual temperature was reported to be 39.62C.

Freq. vs temperature is plotted below in the attached PDF. The slope is 2.8GHz/K.

The data is in the attached MATLAB file.

Attachment 1: LightWave_temp_sweep.pdf
LightWave_temp_sweep.pdf
Attachment 2: LightWave_Temp.m
% plot the data from the Lightwave Temperature sweep

% Lightwave temperature

LWTemp = [0.2744
    0.2753
    .2767
    .2780
    .2794
    .2808
... 67 more lines ...
  2792   Mon Apr 12 17:48:32 2010 AidanUpdateComputer Scripts / Programselog restarted

 The elog crashed when I was uploading a photo just now. I logged into nodus and restarted it.

  2791   Mon Apr 12 17:37:52 2010 josephbUpdateComputersY end simulated plant progress

Currently, the y end plant is yep.mdl.  In order to compile it properly (for the moment at least) requires running the normal makefile, then commenting out the line in the makefile which does the parsing of the mdl, and rerunning after modifying the /cds/advLigo/src/fe/yep/yep.c file.

The modifications to the yep.c file are to change the six lines that look like:

"plant_mux[0] = plant_gndx"  into lines that look like "plant_mux[0] = plant_delayx".  You also have to add initialization of the plant_delayx type variables to zero in the if(feInt) section, near where plant_gndx is set to zero.

This is necessary to get the position feedback within the plant model to work properly.

 

#NOTE by Koji

CAUTION:
This entry means that Makefile was modified not to parse the mdl file.
This affects making any of the models on megatron.

Attachment 1: YEP.png
YEP.png
Attachment 2: YEP_PLANT.png
YEP_PLANT.png
  2790   Mon Apr 12 17:09:30 2010 AlbertoUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Quote:

Quote:
 

 1) True. My bad. In my elog entry (but not in my fit code) I forgot the impedance Z= 750Ohm (as in the fit) of the resonant circuit in front of the square root: Vn = Vdn + Z * sqrt( 2 e Idc )

2) That is exactly the point I was raising! The measured dark noise at resonance is 2x what I expect.

3) I don't have uncertainties for the fit offset (that is, for the Dark Noise). The quick fit that I used (Matlab's Non Linear Least Squares method) doesn't provide 95% confidence bounds when I constrain the offset parameter the way I did (I forced it to be strictly positive).
Sure. It's not a very good fit. I just wanted to see how the data was going.

I also admitted that the data points were few, especially around the 3dB point.

Today I'm going to repeat the measurement with a new setup that lets me tune the light intensity more finely.

 Here's another measurement of the noise of the REFL11 PD.

This time I made the fit constraining the Dark Noise. I realized that it didn't make much sense leaving it as a free coefficient: the dark noise is what it is.

2010-04-09_REFL11NoiseMeasurements.png

Result: the transimpedance of REFL11at 11 MHz is about 4000 Ohm.

Note:
This time, more properly, I refer to the transimpedance as the ratio between Vout @11Mhz / Photocurrent. In past entries I improperly called transimpedance the impedance of the circuit which resonates with the photodiode.
  2789   Mon Apr 12 16:20:05 2010 AlbertoConfiguration40m UpgradingREFL55 improved
During the commissioning of the AS55 PD, I learned how to get a much better rejection of the 11MHz modulation.
So I went back to REFL55 and I modified it using the same strategy. (Basically I added another notch to the circuit).
After a few days of continuous back and forth between modeling, measuring, soldering, tuning I got a much better transfer function.

All the details and data will be included in the wiki page (and so also the results for AS55). Here I just show the comparison of the transfer functions that I measured and that I modeled.

I applied an approximate calibration to the data so that all the measurements would refer to the transfer function of Vout / PD Photocurrent. Here's how they look like. (also the calibration will be explained in the wiki)

2010-04-12_REFL55_TF_model_to_meas_comparison.png.

The ratio between the amplitude of the 55Mhz modulation over the 11MHz is ~ 90dB

The electronics TF doesn't provide a faithful reproduction of the optical response.

  2788   Mon Apr 12 14:20:10 2010 kiwamuUpdateGreen LockingPZT response for the innolight

I measured a jitter modulation caused by injection of a signal into laser PZTs.

The measurement has been done by putting a razor blade in the middle way of the beam path to cut the half of the beam spot, so that a change of intensity at a photodetector represents the spatial jitter of the beam.

However the transfer function looked almost the same as that of amplitude modulation which had been taken by Mott (see the entry).

This means the data is dominated by the amplitude modulation instead of the jitter. So I gave up evaluating the data of the jitter measurement.

  2787   Sun Apr 11 19:05:34 2010 KojiOmnistructureComputersWhere are the laptops?

One dell is in the clean room for the suspension work.

Quote:

I can't find the DELL laptop anywhere in the lab. Does anyone know where it is?

Also one of the two netbooks is missing.

 

  2786   Sun Apr 11 13:51:04 2010 AlbertoOmnistructureComputersWhere are the laptops?

I can't find the DELL laptop anywhere in the lab. Does anyone know where it is?

Also one of the two netbooks is missing.

ELOG V3.1.3-